# Oblack Owhite m #3

### v1

@@ Abstract
Online reinforcement learning (RL) excels in complex, safety-critical domains, yet it faces challenges such as sample inefficiency, training instability, and a lack of interpretability. Data attribution offers a principled way to trace model behavior back to individual training samples. However, in online RL, each training sample not only drives policy updates but also influences future data collection, violating the fixed dataset assumption in existing attribution methods. In this paper, we initiate the study of data attribution for online RL, focusing on the widely used Proximal Policy Optimization (PPO) algorithm. We start by establishing a local attribution framework, interpreting model checkpoints with respect to the records in the recent training buffer. We design two target functions, capturing agent action and cumulative return respectively, and measure each record’s contribution through gradient similarity between its training loss and these targets. We demonstrate the power of this framework through three concrete applications: diagnosis of learning, temporal analysis of behavior formation, and targeted intervention during training. Leveraging this framework, we further propose an algorithm, iterative influence-based filtering (IIF), for online RL training that iteratively performs experience filtering to refine policy updates. Across standard RL benchmarks (classic control, navigation, locomotion) to RLHF for large language models, IIF reduces sample complexity, speeds up training, and achieves higher returns. Overall, these results advance interpretability, efficiency, and effectiveness of online RL.
@@ 1 Introduction
Reinforcement learning (RL) has achieved remarkable success across a wide range of decision-making tasks—from game playing (Mnih et al., 2015; Silver et al., 2016) to robotic control (Andrychowicz et al., 2020), and the alignment of large language models (LLMs) (Ouyang et al., 2022). Among its variants, online RL, which continuously alternates between data collection and policy updates (e.g., A3C (Mnih et al., 2016), PPO (Schulman et al., 2017)), is well-suited to real-time, adaptive, and safety-critical domains such as autonomous driving, as it enables on-the-fly correction of mistakes and rapid adaptation to non-stationary environments (Sallab et al., 2017; Andrychowicz et al., 2020). However, modern online RL faces several challenges, including sample inefficiency, high variance, and training instability, often requiring millions of interactions for convergence and yielding inconsistent performance across runs (Henderson et al., 2018; Yu, 2018; Dulac-Arnold et al., 2019).
These challenges, together with their deployment in high-stakes domains, necessitate a deeper understanding of the operational mechanisms of online RL. To this end, prior work has explored various methods for RL interpretability (Milani et al., 2024; Cheng et al., 2025). While useful, these methods often lack the fine-grained explanations necessary for effective interventions or have limited applicability (see Appendix A for a detailed review of related work). Addressing these limitations requires exploring new paradigms.
In recent years, data attribution (Ilyas et al., 2024) has emerged as a powerful approach for machine learning interpretability, offering a complementary perspective by tracing model behaviors back to training data. This framework further benefits downstream applications such as data selection (Xia et al., 2024), bias mitigation (Wang et al., 2024), fact tracing (Chang et al., 2025), among others. However, applying data attribution to online RL is non-trivial. In online RL, agents continuously interact with their environment; each collected data record not only contributes to policy updates but also influences subsequent data collected by the evolving policy. This violates the core assumptions of traditional data attribution methods, which are designed for static datasets and fixed objectives.
@@
    A principled and flexible framework (Sec. 3). We propose a local data attribution framework for online RL, interpreting model checkpoints w.r.t. the records from the recent training buffer. We define the attribution entity as the atomic unit in PPO training, design two target functions that capture agent actions and cumulative returns, and measure each record’s influence through gradient similarity between its training loss and the target.
@@
    Fresh insights into learning (Sec. 4). We demonstrate the power of our framework through three applications: a) Diagnosis of learning: we show records most harmful for learning feature inaccurate advantage estimates; b) Temporal analysis of behavior formation: we reveal an intriguing phase transition of critical records in shaping agent behaviors; c) Targeted intervention: we show that removing records with the most negative influences can effectively improve model training.
@@
    Improved training (Sec. 5). Building on the targeted intervention, we further develop an iterative influence-based filtering algorithm (IIF) that significantly improves standard online RL training. Across standard RL benchmarks to modern RLHF for large language models, IIF consistently improves sample efficiency, reduces computational cost, and enhances final performance.
@@ 2 Preliminaries
Figure 1: Online reinforcement learning operates in alternating rounds of data collection and policy updates; our local data attribution framework quantifies how individual records from a single round influence different aspects of policy update in that round.
@@ 2.1 Online reinforcement learning
Online RL typically proceeds in alternating training rounds of data collection and model training (Fig. 1). In round $k$, the data collection phase involves the agent executing the current policy $\pi_{\theta^{(k)}}$, sampling experiences over multiple episodes to accumulate $n$ transition records in a rollout buffer $B^{(k)}$. Each record includes $(s_{t},a_{t},r_{t})$ as well as computed quantities, including the action log probability $\log\pi_{\theta^{(k)}}(a_{t}|s_{t})$, estimated value $v_{t}$, and advantage estimate $\hat{A}_{t}$. Model parameters are then updated iteratively starting from $\theta_{0}^{(k)}=\theta^{(k)}$: at optimization step $j$, training on the mini-batch $\mathcal{B}_{j}^{(k)}$ drawn from $B^{(k)}$ updates parameters from $\theta_{j}^{(k)}$ to $\theta_{j+1}^{(k)}$. In this paper, we focus on Proximal Policy Optimization (PPO), a widely used, effective algorithm in various applications (Berner et al., 2019; Andrychowicz et al., 2020; Ouyang et al., 2022).
@@ 3 A Local Data Attribution Framework for Online RL
The key feature of online RL is the circular dependency between data and model—earlier experiences drive policy updates, and updated policies produce new experiences to learn from. The dependency of data on model (red arrows in Fig. 2) is unique to online RL and cannot be addressed by existing attribution methods. Current data attribution methods include
Figure 2: Twofold data influence: driving policy updates, shaping future data generation.
retraining-based (e.g., Ghorbani and Zou (2019)) and gradient-based, with the latter further divided into static and dynamic (Hammoudeh and Lowd, 2024). Retraining-based method are too costly and thus inapplicable. Static methods implicitly assume model parameters are obtained from solving an empirical risk minimization problem over a fixed dataset, which is violated in the non-stationary, sequential data setting here. While dynamic methods (e.g., TracIn) capture the temporal dependencies of training data influences on model parameters, they still fail to account for this key effect of data-model dependency. Technically, one can compute influence scores ignoring this extra channel of influence, but these scores would deviate significantly from the ideal effect we aim to measure, with errors that may compound exponentially over rounds. Furthermore, quantifying influences through this channel is hard: sampling is stochastic and non-differentiable, precluding direct gradient-based analysis.
@@ 3.1 A framework of local data attribution
$f^{\text{return}}(\theta)\coloneqq\mathbb{E}_{(s,a)\sim\pi^{\text{ref}}}\left[\log\pi_{\theta}(a\mid s)\hat{A}^{\text{ref}}(s,a)\right].$
@@
For attribution in round $k$, we set the reference policy $\pi^{\text{ref}}=\pi_{\theta^{(k)}}$, i.e., the policy snapshot at the beginning of the round. This provides a stable and relevant reference that evolves with training progress, which is crucial for meaningful evaluation and thus attribution as the policy improves. Furthermore, since the training rollout buffer $B^{(k)}$ is collected under $\pi_{\theta^{(k)}}$, we can directly use it as the validation dataset. We provide further discussions on this design choice in Sec. 4.3 and Sec. 5.1.
@@
Here, $\nabla_{\theta}f(\theta^{(k)}_{j})$ is the gradient of the target function evaluated at $\theta^{(k)}_{j}$, and $\nabla_{\theta}\mathcal{L}^{\text{PPO}}(\theta^{(k)}_{j},z_{i})$ is the per-sample gradient of the PPO training objective for record $z_{i}$. We also discuss two design choices in Sec. 5.1 which substantially reduce the computational and storage costs of the vanilla TracIn.
@@ 4 Applications of Local Data Attribution
Experimental setup.   We perform evaluation on a diverse suite of RL environments—navigation (FrozenLake and MiniGrid), classic control (Acrobot and LunarLander), driving (Highway), and locomotion (BipedalWalker)—covering discrete and continuous state and action spaces with varying complexity and reward structures. We defer descriptions of environments to Sec. B.1 and PPO training setups to Sec. B.2. Our code is at https://github.com/LDAORL/LDA-ORL.
@@ 4.1 Diagnosis of learning: what features bottom records?
In this section, we analyze the bottleneck hindering learning in online RL. To achieve this, we examine bottom records for $f^{\text{return}}$ and uncover a pattern that universally occurs across training rounds (more examples in Sec. C.1): bottom records feature inaccurate advantage estimates.
@@
Fig. 3(a–b) illustrates two examples. In FrozenLake, bottom records include poor actions receiving high positive $\hat{A}$ and good actions receiving negative $\hat{A}$. Similarly, in MiniGrid, the agent drifts from the goal but receives positive $\hat{A}$. These instances of misleading advantage estimates harm the learning.
We conduct quantitative analysis to characterize what constitutes “inaccurate” advantage estimates. We approximate the true advantage $A^{\pi}(s,a)$ using Monte Carlo (MC) rollouts from each $(s,a)$, averaging over multiple trajectories (details in Sec. C.4). We refer to this as the MC estimate, denoted by $\bar{A}$, and compare it with the advantage estimate $\hat{A}$. We perform analysis in FrozenLake.
Our analysis reveals two key aspects of “inaccuracy”: (1) Sign mismatch: A significant proportion of bottom records exhibit opposite signs for the advantage estimate $\hat{A}$ and the MC estimate $\bar{A}$ (marked by red points in Fig. 3(c)). (2) Large magnitude errors: These records also have large $|\bar{A}-\hat{A}|$. Together, sign flips and large magnitude errors generate strong but misleading learning signals. Indeed, the Spearman rank correlation (Spearman, 1904) between each record’s influence and the product $\bar{A}\cdot\hat{A}$ is strongly negative (Fig. 3(d)), confirming that misaligned advantages drive harmful gradient steps.
@@ 4.2 Temporal analysis of behavior formation: phase transition of top records
We investigate the reinforcement of a specific behavior ($a$ at $s$), characterized by a monotonic increase in $\pi(a|s)$. We track the evolution of top records w.r.t. $f^{\text{action}}$ across training rounds, which are critical in shaping the agent’s behavior. Our analysis reveals an intriguing three-stage phase transition (Fig. 4).
[…]
@@
    Initial association: Initially, top records highlight patterns based on simple action-advantage association: they manifest target action paired with positive $\hat{A}$, or alternative actions paired with negative $\hat{A}$ (see Sec. C.2 for examples). The agent’s behavior in this phase is reinforced through this naive association, largely ignoring the context of state. This basic association persists throughout training, even as more complex relationships are learned.
@@
We quantify these phases by analyzing the roughness (normalized Dirichlet energy) (Von Luxburg, 2007) of a similarity graph, a measure closely related to the graph Laplacian (Chung, 1997). In this graph, nodes represent records, values are ($L_{\infty}$-normalized) influence scores $\tilde{I}_{i}$, edge weights $w_{ij}$ capture semantic similarity and decay with embedding distance (details in Sec. C.2). Roughness, computed as $\nicefrac{{\sum w_{ij}(\tilde{I}_{i}-\tilde{I}_{j})^{2}}}{{\sum w_{ij}}}$, is low when semantically similar records have similar influence; this captures the clustering effect. We track roughness across training rounds. As Fig. 4 shows, roughness remains high in Phase 1, indicating influence scores are largely uncorrelated with semantic similarity. It then significantly drops in Phase 2, representing the formation of semantically meaningful clusters of records with similar influences. In Phase 3, roughness remains low due to the settling of clustering, but exhibits minor fluctuations due to influence scores dominated by noise upon convergence.
@@ 4.3 Targeted interventions during training: filtering amplifies policy gain
Sec. 4.1 demonstrates that our framework can identify harmful training records, thereby opening possibilities for targeted interventions. As a sanity check, we apply a simple intervention procedure within a single training round to verify if removing these records yields performance gains.
Our procedure is straightforward: in round $k$, we identify records in $B^{(k)}$ with negative influence scores w.r.t. $f^{\text{return}}$, remove them, and re-train the agent on the filtered dataset starting from $\theta^{(k)}$. Fig. 5 shows that this consistently improves performance throughout learning and across environments.
@@
Figure 5: Boxplots of $\Delta$ return for single round interventions in two environments; red dashed line for zero $\Delta$. We intervene for each round independently. The $\Delta$ return is computed as the difference between the test return of the model trained on the filtered dataset and the original dataset. Results are shown for $3$ random seeds. Additional results can be found in Sec. C.3.
A reader may ask: how can $f^{\text{return}}$ be meaningful when it relies on on-policy data with potentially inaccurate advantage estimates, unlike clean validation data used in traditional data attribution for supervised learning? Despite potential noise in individual records, the aggregated signal from $f^{\text{return}}$ is reasonably robust. This arises from the close alignment of $f^{\text{return}}$ with the PPO objective: effective PPO updates on the training buffer implies a reliable $f^{\text{return}}$ for attribution, enabling our intervention to clear away misleading records while retaining beneficial ones. This can be seen as purifying the learning signal, thereby amplifying the improvement achieved by PPO. More discussions are in Sec. C.3.

### v2

# A Snapshot of Influence: A Local Data Attribution Framework for Online Reinforcement Learning

@@ Abstract
Online reinforcement learning (RL) excels in complex, safety-critical domains but suffers from sample inefficiency, training instability, and limited interpretability. Data attribution provides a principled way to trace model behavior back to training samples, yet existing methods assume fixed datasets, which is violated in online RL where each experience both updates the policy and shapes future data collection. In this paper, we initiate the study of data attribution for online RL, focusing on the widely used Proximal Policy Optimization (PPO) algorithm. We start by establishing a local attribution framework, interpreting model checkpoints with respect to the records in the recent training buffer. We design two target functions, capturing agent action and cumulative return respectively, and measure each record’s contribution through gradient similarity between its training loss and these targets. We demonstrate the power of this framework through three concrete applications: diagnosis of learning, temporal analysis of behavior formation, and targeted intervention during training. Leveraging this framework, we further propose an algorithm, iterative influence-based filtering (IIF), for online RL training that iteratively performs experience filtering to refine policy updates. Across standard RL benchmarks (classic control, navigation, locomotion) to RLHF for large language models, IIF reduces sample complexity, speeds up training, and achieves higher returns. Together, these results open a new direction for making online RL more interpretable, efficient, and effective.
@@ 1 Introduction
Reinforcement learning (RL) has achieved remarkable success across a wide range of decision-making tasks, from game playing (Mnih et al., 2015; Silver et al., 2016) to robotic control (Andrychowicz et al., 2020) and the alignment of large language models (LLMs) (Ouyang et al., 2022). Among its variants, online RL, which continuously alternates between data collection and policy updates (e.g., A3C (Mnih et al., 2016), PPO (Schulman et al., 2017)), is well-suited to real-time, adaptive, and safety-critical domains such as autonomous driving, as it enables on-the-fly correction of mistakes and rapid adaptation to non-stationary environments (Sallab et al., 2017; Andrychowicz et al., 2020). However, modern online RL faces several challenges, including sample inefficiency, high variance, and training instability, often requiring millions of interactions for convergence and yielding inconsistent performance across runs (Henderson et al., 2018; Yu, 2018; Dulac-Arnold et al., 2019).
These challenges, together with their deployment in high-stakes domains, necessitate a deeper understanding of the operational mechanisms of online RL. To this end, prior work has explored various methods for RL interpretability (Milani et al., 2024; Cheng et al., 2025). While useful, these methods often lack the fine-grained explanations necessary for effective interventions or have limited applicability (see Sec.˜6 for a detailed review of related work). Addressing these limitations requires exploring new paradigms.
In recent years, data attribution (Deng et al., 2025) has emerged as a powerful approach for machine learning interpretability, offering a complementary perspective by tracing model behaviors back to training data. This framework further benefits downstream applications such as data selection (Xia et al., 2024), bias mitigation (Wang et al., 2024), fact tracing (Chang et al., 2025), among others. However, applying data attribution to online RL is non-trivial. In online RL, agents continuously interact with their environment; each collected experience not only contributes to policy updates but also influences future rollouts collected by the evolving policy. This violates the core assumptions of traditional data attribution methods, which are designed for static datasets and fixed objectives.
@@
    A principled and flexible framework (Sec.˜3). We propose a local data attribution framework for online RL, interpreting model checkpoints w.r.t. the records from the recent training buffer. We define the attribution entity as the atomic unit in PPO training, design two target functions that capture agent actions and cumulative returns, and measure each record’s influence through gradient similarity between its training loss and the target.
@@
    Fresh insights into learning (Sec.˜4). We demonstrate the power of our framework through three applications: a) diagnosis of learning: we show records most harmful for learning feature inaccurate advantage estimates; b) temporal analysis of behavior formation: we reveal an intriguing phase transition of critical records in shaping agent behaviors; c) targeted intervention: we show that removing records with the most negative influences can effectively improve model training.
@@
    Improved training (Sec.˜5). Building on the targeted intervention, we further develop an iterative influence-based filtering algorithm (IIF) that significantly improves standard online RL training. Across standard RL benchmarks to modern RLHF for large language models, IIF consistently improves sample efficiency, reduces computational cost, and enhances final performance.
@@ 2 Preliminaries
Figure 1: An Illustration of the alternating learning cycle in online RL (Sec.˜2.1) and our local data attribution framework (Sec.˜3.1). Online RL operates in alternating rounds of data collection and policy updates; our local data attribution framework quantifies how individual records from a single round influence different aspects of policy update in that round.
@@ 2.1 Online reinforcement learning
Online RL typically proceeds in alternating training rounds of data collection and model training (Fig.˜1). In round $k$, the data collection phase involves the agent executing the current policy $\pi_{\theta^{(k)}}$, sampling experiences over multiple episodes to accumulate $n$ transition records in a rollout buffer $B^{(k)}$. Each record contains the raw transition $(s_{t},a_{t},r_{t})$ and several computed quantities, including the action log probability $\log\pi_{\theta^{(k)}}(a_{t}|s_{t})$, estimated value $v_{t}$, and advantage estimate $\hat{A}_{t}$. Model parameters are then updated iteratively starting from $\theta_{0}^{(k)}=\theta^{(k)}$: at optimization step $j$, training on the mini-batch $\mathcal{B}_{j}^{(k)}$ drawn from $B^{(k)}$ updates parameters from $\theta_{j}^{(k)}$ to $\theta_{j+1}^{(k)}$. In this paper, we focus on Proximal Policy Optimization (PPO), a widely used, effective algorithm in various applications (Berner et al., 2019; Andrychowicz et al., 2020; Ouyang et al., 2022).
@@ 3 A Local Data Attribution Framework for Online RL
The key feature of online RL is the circular dependency between data and model—earlier experiences drive policy updates, and updated policies produce new experiences to learn from. The dependency of data on model (red arrows in Fig.˜2) is unique to online RL and cannot be addressed by existing attribution methods. Current data attribution methods include
Figure 2: Twofold data influence: driving policy updates, shaping future data collection.
retraining‑based (e.g., Ghorbani and Zou (2019)) and gradient‑based, with the latter further divided into static and dynamic (Hammoudeh and Lowd, 2024). Retraining-based methods require training the model once for each of the records being evaluated, which is computationally expensive in any setting and particularly prohibitive in RL. Static methods implicitly assume model parameters are obtained from solving an empirical risk minimization problem over a fixed dataset, which is violated in the non-stationary, sequential data setting here. While dynamic methods (e.g., TracIn) capture the temporal dependencies of training data influences on model parameters, they still fail to account for this key effect of data-model dependency. If we compute influence scores using the original formulas from standard supervised learning, they capture only the impact on parameter updates, ignoring the extra channel of influences through future data generation. As a result, the scores may deviate significantly from the true influence we seek to measure. Furthermore, quantifying influences through this channel is challenging because sampling is stochastic and non‑differentiable.
@@ 3.1 A framework of local data attribution
$f^{\text{return}}(\theta)\coloneqq\mathbb{E}_{\tau\sim\pi^{\text{ref}},(s,a)\sim\tau}\left[\log\pi_{\theta}(a\mid s)\hat{A}^{\text{ref}}(s,a)\right].$
@@
For attribution in round $k$, we set the reference policy $\pi^{\text{ref}}=\pi_{\theta^{(k)}}$, i.e., the policy snapshot at the beginning of the round. This is a key design choice of our contextual framework, which enables us to ask: For the agent at its current stage of training, which experiences will be most helpful or harmful for the next update? Unlike a fixed, off-distribution reference that may provide misleading signals due to mismatch with the agent’s current state, our dynamic reference evolves with training, providing a stable and relevant basis for meaningful evaluation and attribution. Furthermore, since the training rollout buffer $B^{(k)}$ is collected under $\pi_{\theta^{(k)}}$, we can directly use it as the validation dataset. We provide further discussions on this design choice in Sec.˜4.3 and Sec.˜5.1.
@@
Remark 1 (Use cases of the two target functions).
The two target functions have different use cases. $f^{\text{action}}$ is mainly for diagnosis: understanding why the agent takes a specific action at a specific state (Sec.˜4.2). On the other hand, $f^{\text{return}}$ assesses contribution to overall performance, which makes it suitable for both analysis (Sec.˜5.2) and algorithmic policy improvement (Sec.˜5).
@@
Here, $\nabla_{\theta}f(\theta^{(k)}_{j})$ is the gradient of the target function evaluated at $\theta^{(k)}_{j}$, and $\nabla_{\theta}\mathcal{L}^{\text{PPO}}(\theta^{(k)}_{j},z_{i})$ is the per-sample gradient of the PPO training objective for record $z_{i}$. We also discuss two design choices in Sec.˜5.1 which substantially reduce the computational and storage costs of the vanilla TracIn.
@@
Remark 2 (Extension to other online RL algorithms).
While we focus on PPO in our study, our framework readily extends to other online RL algorithms. For on-policy methods [^1] For GRPO (Shao et al., 2024), which uses a group-relative baseline rather than value-function baseline, the target function needs to be adjusted as well. such as TRPO (Schulman et al., 2015) and A3C (Mnih et al., 2016), the adaptation only requires modifying the per-sample loss gradient. For offline methods like DQN (Mnih et al., 2013), we need to additionally change the target function to the Bellman error. In all cases, our attribution framework reveals whether training records help or hinder learning at the agent’s current state. A key distinction is that, on-policy methods allow direct validation with current data, whereas off-policy methods require sampling fresh rollouts.
@@ 4 Applications of Local Data Attribution
Experimental setup.   We perform evaluation on a diverse suite of RL environments—navigation (FrozenLake and MiniGrid), classic control (Acrobot and LunarLander), driving (Highway), and locomotion (BipedalWalker)—covering discrete and continuous state and action spaces with varying complexity and reward structures. We defer descriptions of environments to Sec.˜A.1 and PPO training setups to Sec.˜A.2. Our code is at https://github.com/LDAORL/LDA-ORL.
@@ 4.1 Diagnosis of learning: what features bottom records?
In this section, we analyze the bottleneck that hinders learning in online RL. Specifically, we examine the bottom records for $f^{\text{return}}$ and uncover a consistent pattern across training rounds (additional examples in Sec.˜B.1): these bottom records are characterized by inaccurate advantage estimates, echoing observations in the literature (Ilyas et al., 2018).
@@
Fig.˜3(a–b) illustrates two examples. In FrozenLake, bottom records include poor actions receiving high positive $\hat{A}$ and good actions receiving negative $\hat{A}$. Similarly, in MiniGrid, the agent drifts from the goal but receives positive $\hat{A}$. These instances of misleading advantage estimates harm the learning.
We conduct quantitative analysis to characterize what constitutes “inaccurate” advantage estimates. We approximate the true advantage $A^{\pi}(s,a)$ using Monte Carlo (MC) rollouts from each $(s,a)$, averaging over multiple trajectories (details in Sec.˜B.4). We refer to this as the MC estimate, denoted by $\bar{A}$, and compare it with the advantage estimate $\hat{A}$. We perform analysis in FrozenLake.
[…]
@@ 4.2 Temporal analysis of behavior formation: phase transition of top records
We investigate the reinforcement of a specific behavior ($a$ at $s$), characterized by a monotonic increase in $\pi(a|s)$. We track the evolution of top records w.r.t. $f^{\text{action}}$ across training rounds, which are critical in shaping the agent’s behavior. Our analysis reveals an intriguing three-stage phase transition (Fig.˜4).
Figure 4: Phase change of top records in Highway, with the target behavior taking the action “slower” when tailing the front vehicle. In the inner plot, the black curve depicts $\pi(a|s)$; the red curve shows the measured roughness of the graph. Refer to caption: ego vehicle; Refer to caption: other vehicle. Three phases: 1: simple action-advantage associations; 2: semantic clustering (tailing states); 3: no clear patterns.
@@
    Initial association: Initially, top records highlight patterns based on simple action-advantage association: they manifest target action paired with positive $\hat{A}$, or alternative actions paired with negative $\hat{A}$ (see Sec.˜B.2 for examples). The agent’s behavior in this phase is reinforced through this naive association, largely ignoring the context of state. This basic association persists throughout training, even as more complex relationships are learned.
@@
We quantify these phases by analyzing the roughness (normalized Dirichlet energy) (Von Luxburg, 2007) of a similarity graph, a measure closely related to the graph Laplacian (Chung, 1997). In this graph, nodes represent records, values are ($L_{\infty}$-normalized) influence scores $\tilde{I}_{i}$, edge weights $w_{ij}$ capture semantic similarity and decay with embedding distance (details in Sec.˜B.2). Roughness, computed as $\nicefrac{{\sum w_{ij}(\tilde{I}_{i}-\tilde{I}_{j})^{2}}}{{\sum w_{ij}}}$, is low when semantically similar records have similar influence; this captures the clustering effect. We track roughness across training rounds. As Fig.˜4 shows, roughness remains high in Phase 1, indicating influence scores are largely uncorrelated with semantic similarity. It then significantly drops in Phase 2, representing the formation of semantically meaningful clusters of records with similar influences. In Phase 3, roughness remains low due to the settling of clustering, but exhibits minor fluctuations due to influence scores dominated by noise upon convergence.
@@ 4.3 Targeted interventions during training: filtering amplifies policy gain
Sec.˜4.1 demonstrates that our framework can identify harmful training records, thereby opening possibilities for targeted interventions. As a sanity check, we apply a simple intervention procedure within a single training round to verify if removing these records yields performance gains.
Our procedure is straightforward: in round $k$, we identify records in $B^{(k)}$ with negative influence scores w.r.t. $f^{\text{return}}$, remove them, and re-train the agent on the filtered dataset starting from $\theta^{(k)}$. Fig.˜5 shows that this consistently improves performance throughout learning and across environments.
@@
Figure 5: Boxplots of $\Delta$ return for single round interventions in two environments; red dashed line for zero $\Delta$. We intervene for each round independently. The $\Delta$ return is computed as the difference between the test return of the model trained on the filtered dataset and the original dataset. Results are shown for 33 random seeds. Additional results can be found in Sec.˜B.3.
A reader may ask: how can $f^{\text{return}}$ be meaningful when it relies on on-policy data with potentially inaccurate advantage estimates, unlike clean validation data used in traditional data attribution for supervised learning? Despite potential noise in individual records, the aggregated signal from $f^{\text{return}}$ is reasonably robust. This arises from the close alignment of $f^{\text{return}}$ with the PPO objective: effective PPO updates on the training buffer implies a reliable $f^{\text{return}}$ for attribution, enabling our intervention to clear away misleading records while retaining beneficial ones. This can be seen as purifying the learning signal, thereby amplifying the improvement achieved by PPO. More discussions are in Sec.˜B.3.